The following account balances at the beginning of January were selected from the general ledger of Frozen Juice Manufacturing​ Company:


Work in process inventory

​$0

Raw materials inventory

​$28,600

Finished goods inventory

​$41,000

Additional​ data:

1. Actual manufacturing overhead for January amounted to​ $67,200.

2. Total direct labor cost for January was​ $63,100.

3. The predetermined manufacturing overhead rate is based on direct labor cost. The budget for the year called for​ $251,000 of direct labor cost and​ $301,200 of manufacturing overhead costs.

4. The only job unfinished on January 31 was Job No.​ 151, for which total direct labor charges were​ $6,000 (1,700 direct labor​ hours) and total direct material charges were​ $14,600.

5. Cost of direct materials placed in production during January totaled​ $123,600. There were no indirect material requisitions during January.

6. January 31 balance in raw materials inventory was​ $35,000.

7. Finished goods inventory balance on January 31 was​ $34,600.

What is the cost of goods sold for​ January?

A.


​$316,300

B.

​$206,800

C.

​$241,020

D.

​$232,400
Answer #1

Solution:

C. $241,020

Explanation:

Manufacturing overhead % = (301200 / 251000) x 100 = 120%

Unfinished job = 6000 + 14600 + (6000 x 120%) = 27800

Total manufacturing cost = 63100 + 123600 + (63100 x 120%) = 262420

=262420 - 27800 = 234620

Cost of goods sold = 234620 + 41000 - 34600 = 241,020